OpenStudy (anonymous):

What polynomial has roots of –5, –2, and 4 ?

OpenStudy (anonymous):

to do this problem you go like this (x+5) ( x+2) and (x-4)

OpenStudy (anonymous):

x3 + x2 – 22x – 40 ?

OpenStudy (anonymous):

(x+5) * ( x+2) *(x-4) expand by distributive property

OpenStudy (anonymous):

–5, 2, and 4 (x+5)(x-2)(x-4) = (x^2 + 3x - 10)(x-4) = x^3 + 3x^2 - 10x - 4x^2 - 12x + 40 = x^3 - x^2 - 22x + 40

OpenStudy (anonymous):

Awwwww thanks huneya. I was so close!

OpenStudy (anonymous):

x^3 +3*x^2 -18*x -40

OpenStudy (anonymous):

@huneya i think (x+5)(x-2)(x-4) is wrong

OpenStudy (anonymous):

What polynomial has roots of –5, –2, and 4 ? i is (x+5) * ( x+2) *(x-4) is not (x+5)(x-2)(x-4) -5, 2 and -4